OpenStudy (anonymous):

Solve and check. 73/4=X/8 A 1/4 B 62 C 1/62 D 4

OpenStudy (jordanloveangel):

Solution for 73/4=x/8 equation: x in (-oo:+oo) 73/4 = x/8 // - x/8 73/4-(x/8) = 0 73/4-1/8*x = 0 // - 73/4 -1/8*x = -73/4 // : -1/8 x = -73/4/(-1/8) x = 146 x = 146

OpenStudy (anonymous):

So witch is it the